# MARCH: Evaluating the Intersection of Ambiguity Interpretation and Multi-hop Inference

### v2

@@ Abstract
Real-world multi-hop QA is naturally linked with ambiguity, where a single query can trigger multiple reasoning paths that require independent resolution. Since ambiguity can occur at any stage, models must navigate layered uncertainty throughout the entire reasoning chain. Despite its prevalence in real-world user queries, previous benchmarks have primarily focused on single-hop ambiguity, leaving the complex interaction between multi-step inference and layered ambiguity underexplored. In this paper, we introduce MARCH, a benchmark for their intersection, with 2,209 multi-hop ambiguous questions curated via multi-LLM verification and validated by human annotation with strong agreement. Our experiments reveal that even state-of-the-art models struggle with MARCH, confirming that combining ambiguity resolution with multi-step reasoning is a significant challenge. To address this, we propose CLARION, a two-stage agentic framework that explicitly decouples ambiguity planning from evidence-driven reasoning, significantly outperforms existing approaches, and paves the way for robust reasoning systems.
@@ 1 Introduction
Multi-hop Question Answering (QA) presents a significant reasoning challenge, requiring models to construct logical chains by connecting disparate pieces of information scattered across multiple documents (trivedi2022musique; yang2018hotpotqa; ho2020constructing). Ambiguity in QA can further complicate this process, as a single query may stem from polysemous terms or insufficient context, demanding clarification or interpretation before an answer can be derived (min-etal-2020-ambigqa). The intersection of these two challenges—multi-hop reasoning and ambiguity—creates a uniquely difficult setting where uncertainty scales exponentially. In multi-hop ambiguous QA, ambiguity can emerge at any step of the reasoning chain, often remaining latent until prior steps are resolved. This interdependence means that errors in resolving early-stage ambiguity propagate downstream, causing models to prematurely commit to incorrect reasoning paths and producing incomplete or flawed answers.
@@ 2 Multi-Hop Ambiguous QA
We define multi-hop ambiguous QA as a task where a single input query triggers multiple valid reasoning chains, requiring the system to resolve uncertainties that determine the trajectory of multi-step inference. In multi-hop ambiguous QA, each valid interpretation dictates a unique sequence of intermediate decomposition steps (e.g., identifying different bridge entities), which in turn necessitates retrieving disjoint sets of evidence documents. Consequently, failing to resolve ambiguity at the initial or intermediate hops leads to a cascading failure, where the reasoning agent pursues an irrelevant trajectory and cannot recover the correct final answer. To systematically analyze these challenges, we extend the standard ambiguity taxonomy to the multi-hop setting. Following prior ambiguity definitions (tang2025clarifying; tanjim2025disambiguation), we extend the taxonomy to multi-hop QA and group ambiguous questions into three types—semantic, syntactic, and constraint (Table 1).
@@
Constraint ambiguity occurs when an over-specific modifier is unnecessary or mismatched with how evidence is written, causing a valid chain to be pruned early. Example: “What is the capital of the country where the highest mountain in Europe is located?” Many sources disagree on whether the highest mountain in Europe is Mount Elbrus (Caucasus) or Mont Blanc (Alps), and some pages simply say “Europe’s highest mountain” without committing. If the system enforces the modifier literally, it may retrieve only one interpretation and miss the other. A robust strategy is to Generalize (relax) the constraint (e.g., consider both candidates) and then verify the remaining hop (country $\rightarrow$ capital). For more details for this taxonomy, see Appendix I.
@@ 3.1 Dataset Construction
We build MARCH from MuSiQue’s validation set and a subset of its training set. Unlike other multi-hop benchmarks (zhu2024fanoutqa; he2024mintqa) that are narrow in domain or inflate hops with list-style questions (e.g., top-5), MuSiQue enforces connected, dependency-linked reasoning across diverse domains. We first filter out questions from MusiQue that lack ambiguity and retain only those judged as ambiguous by our multi-stage pipeline. Let $\mathcal{Q}_{\text{base}}$ be the set of base multi-hop questions. We consider three ambiguity types $\mathcal{T}=\{\text{Semantic},\text{Syntactic},\text{Constraint}\}$. We use a set of off-the-shelf LLMs as detectors; for a question $q\in\mathcal{Q}_{\text{base}}$, type $t\in\mathcal{T}$, and detector $m$, let $y_{m,t}(q)\in\{0,1\}$ denote whether $q$ is judged ambiguous of type $t$.
@@
For each question of MuSiQue, we provide definitions of each ambiguity type and ask multiple LLMs to detect type-wise ambiguity. We employ four detectors: gpt-4.1 (achiam2023gpt), llama-4-maverick (meta2025llama4maverick), qwen3-235b-a22b (yang2025qwen3), and claude-sonnet-4 (anthropic2025claudesonnet4). We keep a type label only when the detectors are fully in agreement.
@@
To generate answers for the questions, we require evidence obtained via retrieval over clarified questions. However, clarified questions can remain multi-hop, and questioning with a specific multi-hop form may narrow the search scope and miss relevant documents. To mitigate this, we use gpt-4.1 to decompose each clarified question $c\in\mathcal{C}(q,t)$ into atomic sub-questions $\mathcal{S}(c)=\{s_{1},\ldots,s_{k}\}$. For every $s\in\mathcal{S}(c)$, we retrieve up to 10 candidate documents from English Wikipedia [^1] https://dumps.wikimedia.org/. We then pool candidates $\mathcal{D}(c)=\bigcup_{s\in\mathcal{S}(c)}\mathcal{D}(s)$, and if $|\mathcal{D}(c)|<10$, we additionally perform retrieval process with the clarified question $c$ itself to back-fill more evidence. Next, we perform embedding-based re-ranking with Qwen3-8B-Embedding (qwen3embedding) by the similarity between the clarified question and document passages. Finally, we sort $\mathcal{D}(c)$ by this score to prioritize evidence aligned with the clarified interpretation.
@@
Given each clarified question $c$ and its ranked candidate documents $\mathcal{D}(c)$, we use gpt-4.1 to produce a short factual answer only when the retrieved evidence clearly supports it; otherwise, we omit the short answer and drop that clarified item. For retained cases, we also record the passage used for generating the short answer. Finally, for the original question $\mathcal{Q}_{\text{base}}$, we utilize gpt-4.1 to write a single-sentence long answer that connects the two short answers into a coherent statement while incorporating interpretations and citations. For the detailed prompts, refer to Appendix J.
@@
Before filtering, we ensure that short answers remain concise. If either clarified question has a short answer longer than 10 tokens, we cut it down to a much shorter form with gpt-4.1. After that, we remove cases where the two short answers are identical. (We provide representative examples after applying this collision rule in Appendix H.) After the filtering, the final MARCH dataset consists of 2,209 examples. For the final filtering stage, we exclude gpt-4.1—already used in Steps 2 and 3—and instead employ llama-4-maverick, qwen3-235b-a22b, and claude-sonnet-4. Each candidate instance, including the question, clarified questions, ambiguity type, supporting passages, short answers, and long answers, is independently checked for alignment in all fields. We retain only those cases where all three models unanimously judged the instance as fully aligned, using the same criteria as our human evaluation protocol. The upper side of Table 2 shows statistics and reports key characteristics of MARCH. MARCH also retains broad topical coverage; Appendix B reports the domain distribution of MARCH.
@@ 4 CLARION: An Agentic Framework for Multi-hop Ambiguous QA
Model Method STR-EM (%) Disambig-F1 (%) Avg (%) LLM-as-a-Judge
@@
CLARION w/o clarification 25.10 25.56 25.33 2.922
CLARION w/o clarification & detection 22.94 24.02 23.48 2.782
@@
CLARION w/o clarification 24.12 22.54 23.33 2.609
CLARION w/o clarification & detection 23.82 22.04 22.93 2.573
@@
CLARION w/o clarification 23.63 22.99 23.31 2.927
CLARION w/o clarification & detection 24.51 24.31 24.41 2.906
Table 3: MARCH benchmark results with baseline reasoning methods. Metrics are scaled to percentages. Best per-model scores for each metric are in bold, and second-best scores are underlined.
@@
The Acting Agent executes the reasoning plan through a ReAct-style prompting (yao2023react) scheme, unfolding in a Thought $\rightarrow$ Action $\rightarrow$ Observation loop. Unlike naive ReAct, which typically follows a single reasoning trajectory whose early assumptions steer subsequent retrieval, CLARION separates interpretation enumeration from evidence gathering. The Planning Agent produces an initial set of clarified interpretations, and the Acting Agent runs a ReAct loop that retrieves and reasons for each interpretation before synthesizing the final long-form answer. Concretely, the Acting Agent maintains an interpretation set and collects evidence per interpretation, and is instructed to produce a final answer that explicitly covers all interpretations (or as many as possible under the iteration budget). At each iteration, the agent selects one of three actions: (1) Search: retrieve external documents when additional evidence is needed; (2) Planning: refine or expand the current interpretation set (e.g., add a missing interpretation) when the current plan is insufficient; (3) Answer: synthesize the final output once enough evidence has been gathered. To ensure reliable parsing and automated execution, all actions generated by the agent must be in JSON format. Furthermore, to prevent infinite loops and ensure computational tractability, the ReAct prompting is limited to a maximum of five iterations. If this limit is reached without a resolution, the agent is compelled to execute the Answer action, formulating the best possible response based on the information gathered thus far. We provide implementation details(models, decoding, retrieval hyperparameters) and the full prompt templates used at each stage in Appendix H and Appendix J, respectively.
@@ 5 Experiments and Results
We evaluate three state-of-the-art LLMs widely used in real-world scenarios for MARCH: qwen3-235b-a22b-2507 (yang2025qwen3), gemini-2.5-flash (comanici2025gemini), and deepseek-chat-v3.1 (liu2024deepseek). To ensure fair comparison, all systems utilize the same retriever (based on qwen3-embedding-8b) and identical retrieval hyperparameters.
@@
We evaluate performance using three metrics: (1) STR-EM (Strict Exact Match): the percentage of gold short answers that appear in the generated long answer after normalization. (2) Disambig-F1: We use a frozen extractive QA model as an evaluator. Given the model-generated long answer, we treat it as the context and, for each gold clarified question (one per interpretation), the QA model extracts a short answer span. We then compute token-level F1 against the corresponding gold short answer and average over all interpretations. (3) LLM-as-a-Judge: a GPT-4-based judge scores the long answer from 0–5 on Relevance, Faithfulness, Informativeness, and Correctness; we report the average score (validated in Appendix F.1).
@@
Given that multi-hop ambiguous QA inherently requires external knowledge, we employ search-based baselines for all experiments. (1) No Retrieval: LLM-only inference without any external context. (2) CoT (wei2022chain): A standard Chain-Of-Thought prompting without retrieval. (3) NaiveRAG (lewis2020retrieval): A standard retrieve-then-read pipeline that retrieves top-$k$ passages using the original question. (4) CoT with RAG (wei2022chain; lewis2020retrieval): CoT prompting augmented with retrieved documents. (5) DIVA (in-etal-2025-diversify): A diversify–verify–adapt framework designed for ambiguous QA. It diversifies the query into multiple interpretations, verifies evidence for each (labeling passages as Useful/Partially Useful/Useless), and adapts its answering strategy based on evidence sufficiency. (6) ReAct (yao2023react): An agentic baseline that dynamically interleaves reasoning and retrieval. It generates Thoughts to plan, executes retrieval Actions, and uses Observations to iteratively refine its path.
@@
We attribute this failure to the path-dependent nature of multi-hop ambiguity: committing to an early interpretation conditions downstream decomposition and retrieval, which can prevent alternative later-hop meanings from even surfacing. Moreover, evidence for different interpretations is often disjoint (via different bridge entities), so naive top-$k$ retrieval tends to over-focus on the dominant branch and under-cover the others.
@@
AmbigQA 14,042 Ambiguous QA
CondAmbigQA 200 Conditional Ambiguous QA
ASQA 5,301 Long-form QA
AmbigDocs 36,098 Ambiguous QA
Table 5: Comparison of ambiguous QA benchmarks and MARCH.

### v4

@@ Abstract
Real-world multi-hop QA is naturally linked with ambiguity, where a single query can trigger multiple reasoning paths that require independent resolution. Since ambiguity can occur at any stage, models must navigate layered uncertainty throughout the entire reasoning chain. Despite its prevalence in real-world user queries, previous benchmarks have primarily focused on single-hop ambiguity, leaving the complex interaction between multi-step inference and layered ambiguity underexplored. In this paper, we introduce MARCH, a benchmark for their intersection, with 2,209 multi-hop ambiguous questions curated via multi-LLM verification and validated by human annotation with strong agreement. Our experiments reveal that even state-of-the-art models struggle with MARCH, confirming that combining ambiguity resolution with multi-step reasoning is a significant challenge. To address this, we propose CLARION, a two-stage agentic framework that explicitly decouples ambiguity planning from evidence-driven reasoning, significantly outperforms existing approaches, and paves the way for robust reasoning systems. The Code is available at https://github.com/jeonghyunpark2002/MARCH.git
@@ 1 Introduction
Multi-hop Question Answering (QA) presents a significant reasoning challenge, requiring models to construct logical chains by connecting disparate pieces of information scattered across multiple documents (Trivedi et al., 2022; Yang et al., 2018; Ho et al., 2020b). Ambiguity in QA can further complicate this process, as a single query may stem from polysemous terms or insufficient context, demanding clarification or interpretation before an answer can be derived (Min et al., 2020). The intersection of these two challenges—multi-hop reasoning and ambiguity—creates a uniquely difficult setting where uncertainty scales exponentially. In multi-hop ambiguous QA, ambiguity can emerge at any step of the reasoning chain, often remaining latent until prior steps are resolved. This interdependence means that errors in resolving early-stage ambiguity propagate downstream, causing models to prematurely commit to incorrect reasoning paths and producing incomplete or flawed answers.
@@ 2 Multi-Hop Ambiguous QA
We define multi-hop ambiguous QA as a task where a single input query triggers multiple valid reasoning chains, requiring the system to resolve uncertainties that determine the trajectory of multi-step inference. In multi-hop ambiguous QA, each valid interpretation dictates a unique sequence of intermediate decomposition steps (e.g., identifying different bridge entities), which in turn necessitates retrieving disjoint sets of evidence documents. Consequently, failing to resolve ambiguity at the initial or intermediate hops leads to a cascading failure, where the reasoning agent pursues an irrelevant trajectory and cannot recover the correct final answer. To systematically analyze these challenges, we extend the standard ambiguity taxonomy to the multi-hop setting. Following prior ambiguity definitions (Tang et al., 2025; Tanjim et al., 2025), we extend the taxonomy to multi-hop QA and group ambiguous questions into three types—semantic, syntactic, and constraint (Table 1).
@@
Constraint ambiguity occurs when an over-specific modifier is unnecessary or mismatched with how evidence is written, causing a valid chain to be pruned early. Example: “What is the capital of the country where the highest mountain in Europe is located?” Many sources disagree on whether the highest mountain in Europe is Mount Elbrus (Caucasus) or Mont Blanc (Alps), and some pages simply say “Europe’s highest mountain” without committing. If the system enforces the modifier literally, it may retrieve only one interpretation and miss the other. A robust strategy is to Generalize (relax) the constraint (e.g., consider both candidates) and then verify the remaining hop (country $\rightarrow$ capital). For more details for this taxonomy, see Appendix J.
@@ 3.1 Dataset Construction
We build MARCH from MuSiQue’s validation set and a subset of its training set. Unlike other multi-hop benchmarks (Zhu et al., 2024; He et al., 2024) that are narrow in domain or inflate hops with list-style questions (e.g., top-5), MuSiQue enforces connected, dependency-linked reasoning across diverse domains. We first filter out questions from MuSiQue that lack ambiguity and retain only those judged as ambiguous by our multi-stage pipeline. Let $\mathcal{Q}_{\text{base}}$ be the set of base multi-hop questions. We consider three ambiguity types $\mathcal{T}=\{\text{Semantic},\text{Syntactic},\text{Constraint}\}$. We use a set of off-the-shelf LLMs as detectors; for a question $q\in\mathcal{Q}_{\text{base}}$, type $t\in\mathcal{T}$, and detector $m$, let $y_{m,t}(q)\in\{0,1\}$ denote whether $q$ is judged ambiguous of type $t$.
@@
For each question of MuSiQue, we provide definitions of each ambiguity type and ask multiple LLMs to detect type-wise ambiguity. We employ four detectors: gpt-4.1 (Achiam et al., 2023), llama-4-maverick (AI, 2025), qwen3-235b-a22b (Yang et al., 2025), and claude-sonnet-4 (Anthropic, 2025). We keep a type label only when the detectors are fully in agreement.
@@
To generate answers for the questions, we require evidence obtained via retrieval over clarified questions. However, clarified questions can remain multi-hop, and questioning with a specific multi-hop form may narrow the search scope and miss relevant documents. To mitigate this, we use gpt-4.1 to decompose each clarified question $c\in\mathcal{C}(q,t)$ into atomic sub-questions $\mathcal{S}(c)=\{s_{1},\ldots,s_{k}\}$. For every $s\in\mathcal{S}(c)$, we retrieve up to 10 candidate documents from English Wikipedia [^1] https://dumps.wikimedia.org/. We then pool candidates $\mathcal{D}(c)=\bigcup_{s\in\mathcal{S}(c)}\mathcal{D}(s)$, and if $|\mathcal{D}(c)|<10$, we additionally perform retrieval process with the clarified question $c$ itself to back-fill more evidence. Next, we perform embedding-based re-ranking with Qwen3-8B-Embedding (Zhang et al., 2025b) by the similarity between the clarified question and document passages. Finally, we sort $\mathcal{D}(c)$ by this score to prioritize evidence aligned with the clarified interpretation.
@@
Given each clarified question $c$ and its ranked candidate documents $\mathcal{D}(c)$, we use gpt-4.1 to produce a short factual answer only when the retrieved evidence clearly supports it; otherwise, we omit the short answer and drop that clarified item. For retained cases, we also record the passage used for generating the short answer. Finally, for the original question $\mathcal{Q}_{\text{base}}$, we utilize gpt-4.1 to write a single-sentence long answer that connects the two short answers into a coherent statement while incorporating interpretations and citations. For the detailed prompts, refer to Appendix K.
@@
Before filtering, we ensure that short answers remain concise. If either clarified question has a short answer longer than 10 tokens, we cut it down to a much shorter form with gpt-4.1. After that, we remove cases where the two short answers are identical. (We provide representative examples after applying this collision rule in Appendix I.) After the filtering, the final MARCH dataset consists of 2,209 examples. For the final filtering stage, we exclude gpt-4.1—already used in Steps 2 and 3—and instead employ llama-4-maverick, qwen3-235b-a22b, and claude-sonnet-4. Each candidate instance, including the question, clarified questions, ambiguity type, supporting passages, short answers, and long answers, is independently checked for alignment in all fields. We retain only those cases where all three models unanimously judged the instance as fully aligned, using the same criteria as our human evaluation protocol. The upper side of Table 2 shows statistics and reports key characteristics of MARCH. MARCH also retains broad topical coverage; Appendix B reports the domain distribution of MARCH.
@@ 3.2 Dataset Analysis and Validation
Answer Length Statistics.
Table 3 reports token-length percentile statistics for both answer types in March. Each question contains exactly two short answers, one per interpretation, yielding 4,418 short answers in total. We enforced a strict brevity constraint during the filtering stage (Step 4): if any short answer exceeded 10 tokens, we explicitly prompted the LLM to condense it, and discarded instances where this failed. As a result, short answers are highly concise (mean 8.27, median 3.00), confirming that they target precise, extractable spans rather than verbose descriptions. Long answers, which synthesize both interpretations into a coherent statement, are considerably longer (mean 34.86, median 34.00) yet remain focused, with 90% of instances falling within 50 tokens (P90).
Type $n$ Mean Med. P25 P75 P90 P95
Short 4,418 8.27 3.00 2.00 15.00 22.00 27.00
Long 2,209 34.86 34.00 26.00 42.00 50.00 55.60
Table 3: Token-length statistics for short and long answers in March. Each question has exactly two short answers, one per interpretation.
@@ 4 CLARION: An Agentic Framework for Multi-hop Ambiguous QA
Model Method STR-EM Disambig-F1) Avg LLM-Judge
@@
- w/o clarification 25.10 25.56 25.33 2.922
- w/o clarification & detection 22.94 24.02 23.48 2.782
@@
- w/o clarification 24.12 22.54 23.33 2.609
- w/o clarification & detection 23.82 22.04 22.93 2.573
@@
- w/o clarification 23.63 22.99 23.31 2.927
- w/o clarification & detection 24.51 24.31 24.41 2.906
Human 73.00 62.00 67.50
Table 4: MARCH benchmark results. Metrics are scaled to percentages. Bold = best; underline = second best.
@@
The Acting Agent executes the reasoning plan through a ReAct-style prompting (Yao et al., 2023) scheme, unfolding in a Thought $\rightarrow$ Action $\rightarrow$ Observation loop. Unlike naive ReAct, which typically follows a single reasoning trajectory whose early assumptions steer subsequent retrieval, CLARION separates interpretation enumeration from evidence gathering. The Planning Agent produces an initial set of clarified interpretations, and the Acting Agent runs a ReAct loop that retrieves and reasons for each interpretation before synthesizing the final long-form answer. Concretely, the Acting Agent maintains an interpretation set and collects evidence per interpretation, and is instructed to produce a final answer that explicitly covers all interpretations (or as many as possible under the iteration budget). At each iteration, the agent selects one of three actions: (1) Search: retrieve external documents when additional evidence is needed; (2) Planning: refine or expand the current interpretation set when the current plan is insufficient—for instance, when the agent observes evidence inconsistencies such as an inability to ground an interpretation, a broken hop-to-hop connection, or insufficient branch coverage. This Planning action serves as a built-in self-correction mechanism: even if the initial Planning Agent partially misses or misclassifies an ambiguity, the Acting Agent can re-invoke planning to expand or correct the interpretation set and re-initiate retrieval, preventing errors from propagating unchecked through the reasoning chain; (3) Answer: synthesize the final output once enough evidence has been gathered. To ensure reliable parsing and automated execution, all actions generated by the agent must be in JSON format. Furthermore, to prevent infinite loops and ensure computational tractability, the ReAct prompting is limited to a maximum of five iterations. If this limit is reached without a resolution, the agent is compelled to execute the Answer action, formulating the best possible response based on the information gathered thus far. We provide implementation details(models, decoding, retrieval hyperparameters) and the full prompt templates used at each stage in Appendix I and Appendix K, respectively.
@@ 5 Experiments and Results
We evaluate three state-of-the-art LLMs widely used in real-world scenarios for MARCH: qwen3-235b-a22b-2507 (Yang et al., 2025), gemini-2.5-flash (Comanici et al., 2025), and deepseek-chat-v3.1 (Liu et al., 2024). To ensure fair comparison, all systems utilize the same retriever (based on qwen3-embedding-8b) and identical retrieval hyperparameters.
@@
We evaluate performance using three metrics: (1) STR-EM (Strict Exact Match): the percentage of gold short answers that appear in the generated long answer after normalization. (2) Disambig-F1: We use a frozen extractive QA model as an evaluator. Given the model-generated long answer, we treat it as the context and, for each gold clarified question (one per interpretation), the QA model extracts a short answer span. We then compute token-level F1 against the corresponding gold short answer and average over all interpretations. (3) LLM-as-a-Judge: a GPT-4-based judge scores the long answer from 0–5 on Relevance, Faithfulness, Informativeness, and Correctness; we report the average score (validated in Appendix E.1). This evaluation framework mirrors the philosophy of structured, sub-factor evaluation. Rather than assessing the long answer holistically, STR-EM and Disambig-F1 explicitly decompose evaluation into per-interpretation short answers—one for each valid reading of the ambiguous query. This design achieves the same granular diagnostic insight as a JSON-structured evaluation approach, directly revealing which specific semantic branches a model correctly covers and which it misses, without requiring an additional structured gold format.
@@
Given that multi-hop ambiguous QA inherently requires external knowledge, we employ search-based baselines for all experiments. (1) No Retrieval: LLM-only inference without any external context. (2) CoT (Wei et al., 2022): A standard Chain-Of-Thought prompting without retrieval. (3) NaiveRAG (Lewis et al., 2020): A standard retrieve-then-read pipeline that retrieves top-$k$ passages using the original question. (4) CoT with RAG (Wei et al., 2022; Lewis et al., 2020): CoT prompting augmented with retrieved documents. (5) DIVA (In et al., 2025): A diversify–verify–adapt framework designed for ambiguous QA. It diversifies the query into multiple interpretations, verifies evidence for each (labeling passages as Useful/Partially Useful/Useless), and adapts its answering strategy based on evidence sufficiency. (6) ReAct (Yao et al., 2023): An agentic baseline that dynamically interleaves reasoning and retrieval. It generates Thoughts to plan, executes retrieval Actions, and uses Observations to iteratively refine its path.
@@
We attribute this failure to the path-dependent nature of multi-hop ambiguity: committing to an early interpretation conditions downstream decomposition and retrieval, which can prevent alternative later-hop meanings from even surfacing. This error propagation is particularly severe in multi-hop settings because each hop’s output serves as the input to the next—a misresolved ambiguity at hop 1 fixes an incorrect bridge entity, which steers hop 2 retrieval toward an irrelevant trajectory, and so on, compounding errors across the entire reasoning chain.
@@
Human Performance.
To contextualize the difficulty of March, we evaluate human performance on a sampled subset of 60 questions (20 per ambiguity type). Two graduate-level annotators fluent in English answered each question with unrestricted web search access. The annotators achieved an average STR-EM of 73.0 and Disambig-F1 of 62.0, with strong inter-annotator agreement (Cohen’s $\kappa=0.89$). The substantial gap between human performance and the best-performing system (CLARION: STR-EM 38.73, Disambig-F1 28.38) confirms that March poses a significant and unsolved challenge for current reasoning systems, leaving ample room for future progress.
AmbigQA 14,042 Ambiguous QA
CondAmbigQA 200 Conditional Ambiguous QA
ASQA 5,301 Long-form QA
AmbigDocs 36,098 Ambiguous QA
Table 5: Comparison of ambiguous QA benchmarks and MARCH.